# QUIC Delay Control: an implementation of congestion and delay control

### v1

@@ 5.1.2. Multiple Flows
Let us now consider the multi-flow scenario shown in Figure 3, where four connections of the same type share the same channel and are initiated at different start times, i.e. they enter the channel every 10 seconds, and they leave it in a staggered way as well. In this case, the control delay threshold $OWQD_{th}$ is set to 80% for QUIC-DC, and the bottleneck buffer size is set to 2 BDP.
Table 5. Four concurrent flows – emulated link, buffer=2×BDP. (Average $\pm$ standard deviation).
@@
Table 5 presents the average values along with the corresponding standard deviations for the RTT, goodput, Jain’s fairness index, and loss percentage of each control algorithm. As shown, QUIC-DC exhibits the lowest RTT and loss percentage, while maintaining a goodput level comparable to the other considered algorithms. It is worth noting that, since the link capacity is set to 10 Mbps, a goodput around 2.5 Mbps indicates that the link is fairly shared among the four connections.
@@
Figure 6 shows the CDF of the RTT in the case of multiple connections over a shared bottleneck. The figure shows that QUIC-DC provides lower RTT values compared to the other considered congestion control algorithms. Due to space constraints, we are unable to extend our investigation to more complex multiflows scenarios, nor can we explore the friendliness between different types of control algorithms, which are left for further investigation.
@@ 5.2. Real network results
In this section we report results obtained in the real network scenario described in Section 4.
5.2.1. Terrestrial network
During the experiments, we have measured $RTT{min}=30\ ms$, $RTT{max}=60\ ms$, which means that there is a bottleneck buffer size equal to $1\ BDP$. Table 6 shows that QUIC-DC reduces the loss rate by one order of magnitude whereas the goodput is only slightly lower than the other approaches (note that QUIC Westwood+ corresponds to QUIC-DC $(100\%)$). Moreover, Figure 7 shows that QUIC-DC also provides shorter RTTs. It should be noted that algorithms experiencing higher loss rates provide RTT delays that are underestimated because retransmitted packets are not considered in RTT measurements (Karn and Partridge, 1987).
Figure 7. RTT distribution on real terrestrial path.
Table 6. Terrestrial path, single flow.
@@ 5.2.1. Terrestrial network
5.2.2. Starlink LEO network
We consider the satellite scenario depicted in Figure 3. During the experiments, we measured $RTT_{min}=45\ \mathrm{ms}$ and $RTT_{max}\simeq 100\ \mathrm{ms}$, implying a queuing delay of circa 55 ms. This corresponds to a buffer size of approximately $1\times\mathrm{BDP}$. We’ve conducted experiments with a single flow and with multiple concurrent flows with the same methodology shown in Section 5.1.Table 7 shows that BBRv2 obtains a significant larger goodput, a result that is not kept in the case of multi-flows. Table 8 shows that BBRv2 is broken in the multi-flow case, where it gets 1/4 of the QUIC Westwood+ goodput with a loss rate 2.4 times higher.
Figure8 shows the CDF of the RTTs for all the considered algorithms in the case of single flow. QUIC CUBIC obtains a slightly better distribution but with a smaller goodput as shown in Table 7. Figure 9 shows the CDF of the RTTs for all the considered algorithms in the case of multiple flows. In this case QUIC BBRv2 provides the worst distribution even though it provides the worst goodput (only 5 Mbps whereas QUIC Westwood+ provides 19 Mbps). QUIC-DC using different thresholds on the one-way delay can slightly improve the RTT distributions. However, it should also be noted that algorithms experiencing higher loss rates provide RTT delays that are underestimated because retransmitted packets are not considered in RTT measurements (Karn and Partridge, 1987). In other terms, the algorithms with larger loss-rates provide a distribution of RTT that underestimates the effective delay.
Figure 8. RTT distribution on LEO satellite path – single flow
Table 7. Starlink LEO satellite path, single flow.
CCA Gput Tput Loss RTTavg RTTstd
(Mbps) (Mbps) (%) (ms) (ms)
QUIC-DC (20%) 4.81 4.82 0.13 60.58 7.51
QUIC-DC (50%) 18.06 18.08 0.09 62.30 9.41
Westwood+ 20.12 20.29 0.80 63.21 11.82
BBRv2 29.85 30.17 1.04 67.70 27.02
Cubic 15.07 15.19 0.82 59.29 18.31
New Reno 9.45 9.64 1.96 66.05 20.83
Figure 9. RTT distribution on Starlink LEO satellite path – four concurrent flows
Table 8. Starlink LEO satellite path, four concurrent flows.
CCA RTTavg Goodput Loss
(ms) (Mbps, JFI) (%)
QUIC-DC (20%) $66.36\pm 0.35$ $4.05\pm 0.07$ (1.000) $0.10\pm 0.01$
QUIC-DC (50%) $72.29\pm 0.98$ $8.70\pm 0.61$ (0.996) $0.09\pm 0.02$
Westwood+ $81.72\pm 1.27$ $19.17\pm 0.95$ (0.998) $1.02\pm 0.45$
BBRv2 $87.04\pm 10.50$ $5.09\pm 1.60$ (0.931) $2.44\pm 0.63$
Cubic $74.76\pm 2.78$ $15.28\pm 1.63$ (0.992) $0.81\pm 0.09$
New Reno $78.70\pm 2.06$ $14.81\pm 1.49$ (0.992) $0.93\pm 0.41$
@@ 6. Conclusions
In this paper, we presented QUIC-DC, an algorithm that explicitly controls the queuing delay of a communication connection. The core idea is to estimate the connection one-way queuing delay in order to trigger an early reaction to congestion. The idea has been implemented in QUIC using the TCP Westwood+ congestion control. QUIC-DC has been evaluated against QUIC NewReno, QUIC Westwood+, QUIC Cubic, and QUIC BBR2. Experimental results show that the cumulative distribution function of the round-trip time can be shifted to lower values, while the packet loss rate can be reduced by an order of magnitude.

### v3

@@ 5.1.2. Multiple Flows
The multi-flow scenario shown in Figure 3 is considered. In particular, four connections of the same type are started at different times to avoid synchronization. In this case, the control delay threshold $OWQD_{th}=80\%$ and the bottleneck buffer size is set to 2 BDP. We have measured and averaged goodputs, throughputs, losses and RTTs for each connection.
Table 5. Four concurrent flows, emulated link – buffer=2×BDP. (Average $\pm$ standard deviation).
@@
Table 5 presents the average values along with the corresponding standard deviations for the RTT, goodput, Jain’s fairness index, and loss percentage of each control algorithm. As shown, QUIC-DC exhibits the lowest RTT and loss percentage, while maintaining a goodput level comparable to the other considered algorithms. It is worth noting that, since the link capacity is set to 25 Mbps, a goodput around 6 Mbps indicates that the link is fully utilized and fairly shared among the four connectioncs.
@@
Figure 6 shows the CDF of the RTT in the case of multiple connections over a shared bottleneck. It shows that QUIC-DC provides lower RTT values compared to the other considered congestion control algorithms. Due to space constraints, we are here unable to investigate the friendliness between different types of control algorithms.
@@ 5.2. Real network results
We consider the real network scenario described in Section 4. During the experiments, we have measured $RTT{min}=30\ ms$, $RTT{max}=60\ ms$, which means that there is a bottleneck buffer size equal to $1\ BDP$. Table 6 shows that QUIC-DC reduces the loss rate by one order of magnitude whereas the goodput is only slightly lower than the other approaches (note that QUIC Westwood+ corresponds to QUIC-DC $(100\%)$). Moreover, Figure 7 shows that QUIC-DC also provides shorter RTT. However, it should also be noted that algorithms experiencing higher loss rates provide RTT delays that are underestimated because retransmitted packets are not considered in RTT measurements (Karn and Partridge, 1987). In other terms, the algorithms with larger loss-rates provide a distribution of RTT that underestimates the effective delay.
Figure 7. RTT distribution on real network path.
Table 6. Single flow, real network connections.
@@ 6. Conclusions
QUIC DC controls the delay of an Internet connection by measuring the one-way queueing delay to trigger an early reaction to congestion. The idea has been implemented in QUIC using the TCP Westwood+ congestion control and has been evaluated against QUIC NewReno, QUIC Westwood+, QUIC Cubic, and QUIC BBRv2. Experimental results show that the cumulative distribution function of the round-trip time can be shifted to lower values, while the packet loss rate can be reduced by an order of magnitude. BBRv2 has shown serious drawbacks when multiple flows share the bottleneck.
Acknowledgements.
This work has been funded by the “LOREN” Project,
n. 20223Y85JN, under the PRIN (Progetti di Ricerca d’Interesse Nazionale) 2022 programme of the Italian Ministry of Research.